what is the slope of 2x+y=2.
i dont know if i should subtract the two. or get rid of the two next to the x or just leave the x as 0. please help.

Answer by ankor@dixie-net.com(22740)

what is the slope of 2x+y=2.
:
You want to put the equation into the slope/intercept form which is y = mx + b
:
2x + y = 2
subtract 2x from both sides
y = -2x + 2
This is the slope/intercept form, the slope (m) is -2